The length of a rectangular room is 5 yd more than the width. If the area is 300 yd2, find the length and the width of the room.
Answer by addingup(3677)

Let's call width "w", and the length "l" is w+5. Formula for the area is w*l:
w*(w+5) = 300 Multiply on the left:
w^2 + 5w = 300 Subtract 300 from both sides, and I'm going to rearrange the terms:
-300 + 5w + w^2 = 0 Factor and you get:
(-20 + -1w)(15 + -1w) = 0 Now we have two equations, let's solve for each:
-20 + -1w = 0 and 15 + -1w = 0 Solve the first equation:
-20 + -1w = 0 Add 20 to both sides:
-1w = 20 Divide both sides by -1:
w = -20
Now second equation:
15 + -1w = 0 subtract 15, both sides
-1w = -15 Divide both sides by -1:
w = 15
Our answer is w = {-20, 15} Since -20 is negative we cannot use it. Our answer has to be 15 for the width, And the length is 5 more:
Proof: 15 x (15 + 5) = 15 x 20 = 300 Our answer is correct.